OpenStudy (anonymous):

The Fibonacci sequence is defined by F_1 = F_2 = 1 and F_{n + 2} = F_{n + 1} + F_n. Find the remainder when F_{1999} is divided by 5.

OpenStudy (anonymous):

damn i was waiting for an answer a (somewhat) will known result is that every fifth fibonnaci is divisible by 5, but 1999 is one off so not sure what to do in that case

OpenStudy (anonymous):

*well known

OpenStudy (anonymous):

umm...thanks? (though your not even here) XD

OpenStudy (anonymous):

*you're

OpenStudy (anonymous):

maybe there is a pattern for remainders

OpenStudy (anonymous):

maybe :)

OpenStudy (anonymous):

@satellite73 there is indeed; lookup the Pisano periods. my calculus II teacher told me he discovered the same idea back in undergraduate number theory but couldn't prove that the Fibonacci sequence was periodic mod n for any positive integer n

OpenStudy (anonymous):

anyways, we know it's not \(0\) because:$$\gcd(F_n,F_m)=F_{\gcd(n,m)}$$since \(F_5=5\) we have:$$\gcd(F_{1999},F_5)=F_{\gcd(1999,5)}$$now, since \(1999\) is prime, we have \(\gcd(1999,5)=1\) hence \(\gcd(F_{1999},F_5)=F_1=1\) i.e. \(5\) does not divide \(F_{1999}\)

OpenStudy (anonymous):

1,1,2,3,5=0,8=3,13=3,21=1,34=4,55=0,

OpenStudy (anonymous):

hmm not sure what the pattern is

OpenStudy (anonymous):

http://en.wikipedia.org/wiki/Pisano_period#Tables hence the period \(\mod 5\) is \(20\) yielding the cycle \(01123 03314 04432 02241\)

OpenStudy (anonymous):

01123 03314 04432 02241

OpenStudy (anonymous):

thanks wiki! learn something new every day!

OpenStudy (anonymous):

lol :)

OpenStudy (anonymous):

wait...

OpenStudy (anonymous):

oh no

OpenStudy (anonymous):

and thanks @oldrin.bataku for the "pisano period" never saw that before

OpenStudy (anonymous):

but shouldn't it be 1 ?

OpenStudy (anonymous):

sorry, I MEANT to do \(1999\equiv-1\pmod{20}\) hence we look at the last number in the cycle

OpenStudy (anonymous):

i.e. \(F_{1999}\equiv1\pmod5\) yes

OpenStudy (anonymous):

$$01123 03314 04432 0224\color{red}1$$

OpenStudy (anonymous):

that makes sense. Thanks again oldrin :)

OpenStudy (anonymous):

01123 03314 04432 02241 has 20 numbers and 20 goes in to 2000 evenly, count back one and get 1

OpenStudy (anonymous):

good thing you don't have to prove this i guess

OpenStudy (anonymous):

np @orple8 this one I kept forgetting how to solve... I've done a proof before demonstrating periodicity for said teacher but I'm not sure how to prove \(\pi(5)=20\)

OpenStudy (anonymous):

my guess is induction since almost anything with fibonacci means induction

OpenStudy (anonymous):

I did it using a little bit of group theory I think... it was a semester or so ago

OpenStudy (anonymous):

I checked it using by mathematica, and it is 1

OpenStudy (anonymous):

or just: http://www.wolframalpha.com/input/?i=F_1999+mod+5

OpenStudy (anonymous):

true :)

OpenStudy (anonymous):

\[F(n)= \frac{(\frac{1+\sqrt5}{2})^n-(\frac{1-\sqrt5}{2})^n}{\sqrt5}\]

OpenStudy (anonymous):

yes @cinar that is Binet's formula. you can derive it trivially trying to solve the IVP \(F_n=F_{n-1}+F_{n-2},F_1=F_2=1\) with the ansatz \(F_n=r^n\)

OpenStudy (anonymous):

for example:$$r^n=r^{n-1}+r^{n-2}\\r^n-r^{n-1}-r^{n-2}=0\\r^{n-2}(r^2-r-1)=0$$hence we see that \(r=0\) yields the trivial solution \(F_n=0\) but that the solutions to \(r^2-r-1=0\) give a significantly more interesting one...$$r^2-r-1=0\\r=\frac{1\pm\sqrt{1^2-4(-1)}}2=\frac{1\pm\sqrt5}2$$hence we get \(r=\varphi,1/\varphi+1\) i.e. the golden ratio and its conjugate

OpenStudy (anonymous):

oops that should be \(r=\varphi,1-\varphi\) thus we get two 'fundamental' solutions \(F_n=\varphi^n,F_n=(1-\varphi)^n\) as candidates and (just like linear differential equations) we find that any linear combination of these two satisfies our equations similarly i.e. \(F_n=A\varphi^n+B(1-\varphi)^n\). consider our initial conditions, \(F_1=F_2=1\):$$F_1=1\\A\varphi+B(1-\varphi)=1\\A\varphi-B\varphi+B=1\\(A-B)\varphi=1-B\\F_2=1\\~\\A\varphi^2+B(1-\varphi)^2=1\\A\varphi^2+B(1-2\varphi+\varphi^2)=1\\A\varphi^2+B\varphi^2-2B\varphi+B=1\\(A+B)\varphi^2-2B\varphi=1-B$$using the fact \(\varphi^2=1+\varphi\):$$(A+B)(1+\varphi)-2B\varphi=1-B\\A+B+(A-B)\varphi=1-B$$equating both equations of \(1-B\):$$A+B+(A-B)\varphi=(A-B)\varphi\\A+B=0\\B=-A$$hence:$$A-A+(A+A)\varphi=1+A\\2A\varphi=1+A\\2A\varphi-A=1\\A(2\varphi-1)=1\\A=\frac1{2\varphi-1}=\frac1{\sqrt5}$$ hence \(B=-\dfrac1{\sqrt5}\)

OpenStudy (anonymous):

hence:$$F_n=\frac1{\sqrt5}(\varphi^n-(1-\varphi)^n)$$